OpenStudy (anonymous):

A stone was dropped off a cliff and hit the ground with a speed of 120 ft/s. What was the height of the cliff?

OpenStudy (lgbasallote):

use the formula \[V_2^2 = V_1^2 + 2ah\] where V_1 is initial speed V_2 is final speed a is acceleration h is height since it was dropped, initial velocity is 0 final speed is 120 acceleration is 9.8 (because of gravity) so just substitute \[\implies (120)^2 = 0^2 + 2(9.8) h\] solve for h does that help?

OpenStudy (anonymous):

Thats a physics based equation right? this is in my calc book and the teacher called me out once on using physics making me use calc to do it. I know it sounds crazy but any ideas?

OpenStudy (lgbasallote):

i see. interesting

OpenStudy (anonymous):

yea trust me I want to do it with projectial motion equation myself but she is not taking it

OpenStudy (lgbasallote):

well..we know that the derivative of distance is speed...and the derivative of speed is acceleration are you studying differential or integral calculus?

OpenStudy (anonymous):

Integral, so does that mean I take the double integral of 120?

OpenStudy (lgbasallote):

if you take the double integral of that you get 0 lol

OpenStudy (lgbasallote):

you need a function...

OpenStudy (anonymous):

alright how do we find that?

OpenStudy (lgbasallote):

ugh i just wrote something and it got lost =_=

OpenStudy (anonymous):

oh man :(

OpenStudy (lgbasallote):

i tried taking the double integral of that equation above. do you think it'll work?

OpenStudy (anonymous):

wouldn't that make it 120 -> 120x -> 60x^2 making it the normal tejectory of a projectile?

OpenStudy (anonymous):

and would it let me solve for the height?

OpenStudy (anonymous):

no it would let me solve for time and i wouldn't be able to figure out the height with out physics 8(

OpenStudy (anonymous):

once you know the time you can solve for distance

OpenStudy (lgbasallote):

hmm you make a good point about that integral..i was thinking derivatives

OpenStudy (lgbasallote):

so what would x be?

OpenStudy (anonymous):

x would be time

OpenStudy (lgbasallote):

but you dont have time...

OpenStudy (anonymous):

i should have put a t

OpenStudy (anonymous):

\[\int\limits_{0}^{120} dv=\int\limits_{0}^{t _{f}} adt\] \[120-0=a*(t _{f}-0)\] \[\int\limits dv = \int\limits a*dt= v = a*t + c\]

OpenStudy (anonymous):

\[\int\limits dx =\int\limits v*dt\] \[\int\limits\limits dx =\int\limits\limits (a*t+c)*dt\] \[x =(a*t^2+c*t+c _{2})\]

OpenStudy (anonymous):

awesome man I think i got it from here

OpenStudy (anonymous):

you figured out how to solve for the constants?

OpenStudy (anonymous):

no i thought i did for a sec but it didn't work out

OpenStudy (anonymous):

v = 0 at t=0 x=0 at t=0 so all of our constants are zero

OpenStudy (anonymous):

i screwed up earlier \[x=(\frac{1}{2}*a*t)\]

OpenStudy (anonymous):

oh man i'm wicked confused now

OpenStudy (anonymous):

ok, let's back up and solve for our constants of integration using v =0 at t=0, and x=0 when t=0

OpenStudy (anonymous):

we solved for this earlier \[v=at+c\]

OpenStudy (anonymous):

now we can substitute our initial condition that v=0 when t=0 (0)=a*(0)+c

OpenStudy (anonymous):

0=0+c c=0

OpenStudy (anonymous):

so v=at

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yup that sounds good